### v2

# Supermassive black hole growth from stellar binary encounters

Abstract
The growth of supermassive black holes (SMBHs) remains a central problem in astrophysics, with current observations providing only limited constraints on the underlying mechanisms. One possible growth channel is stellar accretion via the Hill’s mechanism, wherein a SMBH tidally breaks up a passing binary star, capturing and eventually accreting a member of the binary. We adopt a model to predict capture rates from parameters that include the central number density of stars, the stellar velocity dispersion, the binary fraction, and black hole mass. We then use this model to estimate the growth of SMBHs across a range of galactic environments. In a data set of over 90 galaxies of various types and masses, we identify two candidates with SMBHs for which stellar accretion may be a driver of recent growth. Closer to home, a recent analysis of observed hypervelocity stars from the Large Magellanic Cloud (LMC) implicates binary star interactions with a massive black hole. Every hypervelocity star produced in this way leaves a bound partner that may be accreted, providing an active growth channel for the LMC’s black hole.
keywords:
Supermassive black holes (1663) — stellar dynamics (1596) — tidal disruption (1696)
1 Introduction
Enshrouding millions to billions of solar masses within an event horizon of remarkably small size, supermassive black holes (SMBHs) are famously elusive. Yet observations have teased out irrefutable evidence for their existence. In our own galaxy, years-long tracking of stars on tight orbits about a common point beautifully reveal Sgr A, the Milky Way’s four million solar-mass black hole (Ghez et al., 2008; Genzel et al., 2010). The Event Horizon Telescope has since mapped gas flows near the supermassive black hole, unveiling a “shadow” surrounding the event horizon (Event Horizon Telescope Collaboration et al., 2022).
Before these achievements were technologically feasible, Hills (1988) sought other, circumstantial evidence for supermassive black holes. Focusing on the Milky Way, Hills calculated that a binary star passing close to a SMBH in the center of the Galaxy woud be broken up, with one stellar component flung out at high speed in a gravitational sling-shot, while its partner would remain bound to the black hole. The unbound star would be traveling fast enough to escape the Galaxy. Detection of a hypervelocity star (HVS) produced this binary disruption (“break-up”) mechanism would be compelling evidence of the black hole itself.
The discovery of a hypervelocity star by Brown et al. (2005) confirmed Hills’ prediction that a supermassive black hole in the center of the Milky Way produces fast stars with no other plausible origin. Subsequent detections (e.g., Brown et al., 2006, 2009; Zheng et al., 2014, see also Verberne et al. 2025) built up a census of HVSs. The HVS population thus contributed to the growing body of evidence demonstrating that Sgr A was indeed a SMBH. More recently, and based on astrometry from Gaia (Gaia Collaboration et al., 2023), some of these same stars now appear to originate from the Large Magellanic Cloud (LMC). Han et al. (2025) analyzed these sources to infer the existence of LMC, a $\sim{6}\times 10^{5}$ $\rm M_{\odot}$ black hole in the dwarf galaxy. The Hills mechanism is fulfilling its promise as a signpost for SMBHs.
Hypervelocity stars are the lead astronomical actors in the Hills mechanism, yet they are only half the story. The bound partner in each binary break-up remains close to the black hole, and eventually may be accreted onto it (e.g., Bromley et al., 2012). A signature of this process is a tidal disruption event (TDE), which occurs when a black hole’s tidal forces shred the star (Hills, 1975; Frank & Rees, 1976; Lightman & Shapiro, 1977; Rees, 1988, see also Gezari 2021 and references therein). This outcome is possible for black holes with masses less than $\sim{10}^{8}$ $\rm M_{\odot}$, for which the typical tidal disruption radius is outside the event horizon. With more massive black holes, the star simply strays across the event horizon without a trace.
Any bound star can be consumed, either tidally disrupted or swallowed by a black hole, once its orbital angular momentum drifts below a critical value, allowing it to plunge toward the event horizon (e.g., Merritt, 2013; Chen et al., 2023). As fuel for SMBHs, higher-mass stars are more susceptible to tidal disruption because of their diffuse outer layers, which experience weaker gravitational binding forces. As a result, when these stars approach an SMBH, their outer envelopes are easily stripped away by tidal forces (Chen et al., 2023). These partial disruptions can occur multiple times before a massive star is completely destroyed (Ryu et al., 2022). In any case, these stars, as well as their lower-mass counterparts, may ultimately be accreted by the SMBH (e.g., Thomsen et al., 2022; Webb & Foustoul, 2022; Olejak et al., 2025). Stellar accretion may indeed be a dominant early-stage growth channel for black holes in the mass range of $10^{2}$$10^{6}$ $\rm M_{\odot}$, including single-star captures (Stone et al., 2017) and binary break-ups (González Prieto et al., 2025).
A variety of mechanisms have been proposed to explain the growth of supermassive black holes, including mergers and gas accretion (e.g., Silk & Rees, 1998; Yu & Tremaine, 2002; Di Matteo et al., 2003; Montesinos Armijo & de Freitas Pacheco, 2011; Ryu et al., 2022). Yet no single channel has emerged to account for the evolution of all SMBHs. Here, our focus lies on the role of binary stars and the Hills mechanism as contributors to SMBH growth. We begin in §2 by revisiting the theoretical description of this growth channel from Bromley et al. (2012), implementing it in a Python code to track binary star interaction rates with a SMBH and the subsequent evolution of the black hole mass. In §3, we apply the model to galactic data provided by Hannah et al. (2024), which span a wide range of black hole masses and galactic host properties. One feature of this analysis is a prediction for the relationship between black hole mass and central stellar densities, both of which are critical to the rate of disruption of stellar binaries. Closer to home, we assess the growth of LMC, the black hole candidate in the Large Magellanic Cloud S4. We summarize our results in §5.
2 The model: black hole growth through the Hills mechanism
The Hills mechanism operates when the center of mass of a binary star comes within a distance $r_{\rm close}$ of a SMBH, where tidal forces from black hole overcome the gravitational binding of the binary pair (Hills, 1988). If $r_{\rm close}$ is within the event horizon of the SMBH (for black holes with mass $_{\bullet}\gtrsim 10^{8} M\rm M_{\odot}$), or inside the tidal radius of an individual star (for smaller black holes), the binary is certainly destroyed.
At larger closest-approach distances, tidal forces from the black hole are weaker. The gravitational influence of the binary at $r_{\rm close}$ is characterized by
$r_{\rm D}\equiv{r_{\rm close}}\left[\frac{m_{\rm bin}}{11M_{\bullet}}\right]^{1/3}$ (1)
where $m_{\rm bin}$ is the total binary mass (Hills, 1988). If this measure is greater than the binary’s orbital separation $a_{\rm bin}$, then the binary generally will survive its encounter with the SMBH. [^1] The quantity $r_{\rm D}$ is related to the Hill radius, $r_{\rm H}\approx 1.54r_{\rm D}$, from planetary dynamics, describing the gravitational influence of a smaller body on a circular orbit around a larger one. Our choice is also similar to J. Hills’ parameter $D$, which is approximately $175r_{\rm D}/a_{\rm bin}$. For example, a stellar binary with total mass of 2 $\rm M_{\odot}$and a semimajor axis of 1 au can safely pass a $10^{6}$ $\rm M_{\odot}$ black hole if $r_{\rm close}$ is greater than 175 au.
At intermediate closest approach distances, between stellar disruption and $r_{\rm D}$, outcomes depend on details of the orbits including binary separation, orientation, and orbital phase, as well the masses of the two stars and the black hole. A key to Hills’ approach is that these details are treated probabilistically. On the basis of numerical trials with randomized orbital configurations, Hills quantified the probability that a binary encounter with a SMBH will lead to binary break-up entirely in terms of $r_{\rm D}$ and the binary’s orbital separation:
$P_{\text{bbu}}\approx 1-\frac{r_{\rm D}}{a_{\rm bin}}.$ (2)
Distributions of binary properties in a stellar population can inform the range of break-up outcomes, and these outcomes can further be translated into likely HVS ejection speeds and captured star orbital elements based on energy and angular momentum conservation (Hills, 1988; Bromley et al., 2006).
2.1 Binary break-up rates
Here we are interested in average or representative binary properties to predict the change in mass of a SMBH from the accretion of stars captured by the Hills mechanism. Our starting point for the rate is $nv\sigma$” from kinematics, where $n$ is the number density of binaries, $v$ is their mean speed in a region surrounding a SMBH, and $\sigma$ is the cross-sectional area of the black hole — the size of the hypothetical target for a binary star to hit in order to be split into a hypervelocity star and a bound partner. We consider each factor in turn.
We express the number density of binary stars in the vicinity of a SMBH, $n_{\rm bin}$, in terms of the central stellar mass density of its galactic host, $\rho_{0}$, treated as a constant. If the fraction of binaries in this region is $f_{\rm bin}$, then
$n_{\rm bin}\approx\frac{\rho_{0}}{m_{\rm avg}}\frac{f_{\rm bin}}{(1+f_{\rm bin})},$ (3)
where $m_{\rm avg}\approx$ 0.3 $\rm M_{\odot}$ is the average mass of a star, counting binary components as individual stars. In the Milky Way, the binary fraction overall is roughly half of all stars (Duchêne & Kraus, 2013, and references therein), though it is dependent on spectral type of the components (e.g., Moe & Di Stefano, 2017). Typically younger, more massive stars have a higher binary fraction than less massive, older stars.
The central density of stars, $\rho_{0}$, depends strongly on the SMBH mass, $M_{\bullet}$. For large elliptical galaxies, the density falls with increasing black hole mass $\rho_{0}\sim 1/M_{\bullet}$ (e.g., Faber et al., 1997). The trend is opposite in smaller galaxies. Both central stellar density and black hole mass increase with the size and mass of the galactic host. Here, we introduce the parameterization,
$\rho_{0}=5\times 10^{4}\frac{M_{\bullet}/M_{b}}{1+M_{\bullet}^{2}/M_{b}^{2}}\ \ \text{$\rm M_{\odot}$/pc${}^{3}$},$ (4)
where $\rho_{0}$ is the central density within some small (e.g., 5 pc) of the SMBH, and $_{b}\approx 4\times 10^{7} M\rm M_{\odot}$  is the black hole mass associated with the break between two power-law relationships; $\rho_{0}\sim M_{\bullet}$ in the low-mass regime and $\rho_{0}\sim 1/M_{\bullet}$ for masses much greater than $M_{b}$. In the next section, We demonstrate the effectiveness of Equation (4) using data kindly provided by Hannah et al. (2024) that inspired our choice of parameterization.
The central velocity dispersion of stars in the cores of galaxies, $v_{\rm disp}$, is typically of the order of 100 km/s, with values that increase with SMBH mass and total galactic mass (e.g., Faber et al., 1997; Bernardi et al., 2003; McConnell & Ma, 2013; Greene et al., 2020). The range of observed values is narrow compared with that of the masses of known SMBHs, because $v_{\rm disp}$ trends as the fourth root of $M_{\bullet}$ (Kormendy & Ho, 2013) and broader galactic properties including luminosity (e.g., Faber & Jackson, 1976), both of which can span many orders of magnitude.
The third factor in the kinematic description is $\sigma_{\bullet}$, the cross-sectional area of the black hole in scattering events with binaries. This term depends on uncertain quantities including the distribution of binary orbital distances/semimajor exes and the nature of their orbits around the black hole. Following our review of the Hill mechanism above, we adopt $r_{\rm close}$ corresponding to values where the break-up $P_{\text{bbu}}\gtrsim 0.5$ as a measure of the cross-section near the black hole.
The strongest impact of black hole mass is the boost to its cross-sectional area by the long-range effect of gravity on approaching stars. The aim of a binary toward the black hole may be poor, yet gravity draws it in. The cross section is increased by a gravitational focusing factor
$f_{\text{grav}}\approx 1+\frac{2GM_{\bullet}}{r_{\rm close}v_{\rm disp}^{2}},$ (5)
(Bondi & Hoyle, 1944).
Putting these factors together to derive an $nv\sigma$” interaction rate, we find that the break-up rate of binary stars through the Hills mechanism is
$k$ $=$ $\frac{\rho_{0}}{m_{\rm avg}}\frac{f_{\rm bin}}{(1+f_{\rm bin})}\cdot v_{\rm disp}\cdot\pi f_{\text{grav}}r_{\rm close}^{2}$ (6)
$\approx$ $0.0011\left[\frac{\rho_{0}}{10^{4}\ \text{\footnotesize$\rm M_{\odot}$/pc}^{3}}\right]\left[\frac{m_{\rm avg}}{0.3\ \text{\footnotesize$\rm M_{\odot}$}}\right]^{-1}\left[\frac{f_{\rm bin}}{0.5}\right]\times$ (7)
$\left[\frac{v_{\rm disp}}{100~\text{\footnotesize km/s}}\right]^{-1}\left[\frac{M_{\bullet}}{10^{7}~\text{\footnotesize$\rm M_{\odot}$}}\right]\left[\frac{r_{\rm close}}{200~\text{\footnotesize au}}\right]\ \text{yr}^{-1};$ (8)
in the lower expression, we have simplified the dependence on $f_{\rm bin}$ (formally valid for $f_{\rm bin}\ll 1$) and the gravitational focusing factor (since $f_{\text{grav}}\sim M_{\bullet}$ when $f_{\text{grav}}\gg 1$, which holds for all SMBHs and stars in their proximity). This value represents an upper bound, as the fiducial central density $\rho_{0}$ is near its maximum.
While the explicit dependence on parameters is instructive, and indeed is our choice in calculations of black hole growth (below), we also provide an expression with the full dependence on black hole mass, incorporating Equation (4), the $M_{\bullet}$-$v_{\rm disp}$ relation (Kormendy & Ho, 2013), and the potential variation of $r_{\rm close}$ with $M_{\bullet}$ (Hills, 1988):
$k\sim\frac{M_{\bullet}^{25/12}}{M_{b}^{2}+M_{\bullet}^{2}}\rightarrow\begin{cases}M_{\bullet}^{2.08}&\text{if }M_{\bullet}\ll M_{b},\\ M_{\bullet}^{0.08}&\text{if }M_{\bullet}\gg M_{b}.\end{cases}$ (9)
This equation suggests a “sweet spot” for binary break-up around $M_{\bullet}\approx M_{b}$, around $10^{7}$ $\rm M_{\odot}$. Next we explore how these results translate into black hole growth.
2.2 A channel for black hole growth
Every interaction between a binary star and a supermassive black hole that leads to binary break-up places a star on a bound orbit that will eventually lead to tidal disruption as a result of orbital diffusion (e.g., Bromley et al., 2012). While tidal disruption outside the horizon may not result in accretion of all the freed stellar gas, at least initially (Evans & Kochanek, 1989), we assume that for stars already on bound orbits, all the disrupted gas ultimately is accreted by the SMBH. In this scenario, the binary break-up rate in Equation (6) directly provides the black hole growth rate,
$\frac{dM_{\bullet}}{dt}=km_{\rm avg},$ (10)
from accretion of captured Hills’ stars. To solve this equation, we make the following definitions and assumptions:
  • The black hole mass $M_{\bullet}$ is a function of time $t$, with a initial mass $M_{\bullet}(0)$ at $t=0$, the present epoch.
  • The central stellar density $\rho_{0}$ is also a function of time through its dependence on $M_{\bullet}(t)$, wherein $\rho_{0}\sim M_{\bullet}/(M_{b}^{2}+M_{\bullet}^{2})$ (cf. Eq. (4)), scaled so that $\rho_{0}(0)$ is the present-day value.
  • The gravitational focusing factor $f_{\text{grav}}$ depends on $M_{\bullet}(t)$, with $f_{\text{grav}}\sim M_{\bullet}$, since $f_{\text{grav}}\gg 1$ for SMBHs in the cores of galaxies (see Eq. (5)).
  • All other parameters ($r_{\rm close}$, $v_{\rm disp}$, $f_{\rm bin}$) are set to representative constant values, since taken together, they have weak dependence on the SMBH mass.
With this prescription in the growth equation (Eq. (10)), we obtain an ordinary differential equation of the form
$\frac{dM_{\bullet}}{dt}=A\frac{M_{\bullet}^{2}}{M_{\bullet}^{2}+M_{b}^{2}},$ (11)
where $A$ is a constant (cf. Eq. (7)). Its solution, from separation of variables, is
$M_{\bullet}(t)=\frac{1}{2}\left(X+\sqrt{X^{2}+4M_{b}^{2}}\right)$ (12)
where $X=At+M_{\bullet}(0)-M_{b}^{2}/M_{\bullet}(0)$ and $M_{\bullet}(0)$ is the black hole mass at $t=0$, the present epoch.
Here, we also use an explicit, step-wise numerical integrator (the Euler method), written in Python, to solve for black hole growth. We start with the initial SMBH mass $M_{\bullet}(0)$, evaluate all parameters ($\rho_{0}$, $f_{\text{grav}}$, $v_{\rm disp}$, etc.) at that time, and calculate $km_{\rm avg}$. The mass at a time $t=\Delta t$ later is $M_{\bullet}(0)+km_{\rm avg}\Delta t$. This process is repeated to step through to some desired ending time. In practice we use $\Delta t=0.1$–1 Myr. This numerical approach allows us to explore more complicated parameterizations and growth scenarios. Figure 1 provides an illustration.
Figure 1: The evolution of black hole mass from the Hills mechanism and stellar accretion. Each curve shows a unique starting mass, which we evolve usng the growth equation (Eq. (10)). We adopt the analytical relationship between black hole mass and central stellar density in Equation (4), while all other parameters are assigned their fiducial values as in Equation (7). Because the central density is near a maximum for $_{\bullet}\approx 4\times 10^{7} M\rm M_{\odot}$, the most rapid growth occurs for black holes around this size.
3 SMBH growth in nearby galactic hosts
We apply our binary break-up model to observed galactic sources to evaluate the potential growth of supermassive black holes from binary break-up and stellar accretion under varying conditions. While many of the parameters of our model are unknown, estimates are available for the two critical ones, SMBH mass $M_{\bullet}$ and the central stellar density $\rho_{0}$. Here, we consider black hole growth using a curated dataset from Hannah et al. (2024), consisting of 91 galaxies of a range of types, all within 50 Mpc. With the Python Pandas package, we selected data by filtering to include only black holes with mass derivation methods corresponding to source codes 1, 2, and 3, ensuring methodological consistency. We thus identified 30 well-measured galaxies to use in this study. The remaining 61 sources have black hole masses that are not well constrained. For each selected galaxy, we extracted the galaxy name, the black hole mass (in logarithmic solar mass units) and central stellar density (in logarithmic solar masses per cubic parsec), defined as an average inside a distance of 5 pc from the dynamical center of the galaxy.
Figure 2 illustrates the dataset, showing all of the sources in Hannah et al. (2024), including measurements not considered in this study. The plot also shows our analytical expression that broadly tracks the central density as a function of black hole mass (Eq. (4)).
Figure 2: Central density versus SMBH mass. The larger, violet points correspond to sources with more reliable black hole mass estimates, while the smaller, turquoise points represent galaxies with poorly constrained black hole masses. The gray line is a simple analytical model for the mass-density relation (Eq. (4)). The low-mass end of this model is speculative.
To assess the potential for growth, we start with the estimated SMBH mass of each galaxy and the reported central stellar density. We then integrate the mass forward in time as in Figure 1, with an important difference: The central stellar density $\rho_{0}$ is given the same form of dependence on black hole mass as in Equation (4), but the normalization is set to match the density in Hannah et al. (2024). Our results are presented in Figure 3, which shows predicted future growth of the 30 galaxies that we selected from the Hannah et al. (2024) dataset.
Figure 3: Black hole mass and central density evolution from the Hills mechanism growth channel predicted for the 30 best-measured galaxies with SMBHs in the Hannah et al. (2024) sample. We begin with the observed masses and central densities, the use these values to predict future growth of the SMBH, as described in the main text (§2.2). The colors run from green hues (low-mass SMBHs) through reds (intermediate massed) to blue (high-mass SMBHs), and the same unique color is assigned to each galaxy. The thickness of the lines correlates with the relative increase in black hole mass over the 5 Gyr run. Clearly, growth favors the black hole in a high-density field of stars.
Another view of black hole growth comes from integrating the growth equation (Eq. (10)) backward in time to assess the fraction of the observed black hole mass may have arisen from binary break-ups. In this way, we predict the fraction of a black hole’s mass that could potentially have come from the Hills mechanism and stellar accretion. Figure 4 provides that view.
Figure 4: Fraction of the black hole mass predicted to have come from stellar accretion through the Hills mechanism. The diamond symbols are the predictions for the 30 galaxies with high-quality SMBH masses in the (Hannah et al., 2024), run backwards in time for 1 Gyr with our numerical integrator. The short vertical lines near the symbols are the predicted fraction under the assumption that the central density as been fixed in time. For black holes smaller than $_{b}\sim 4\times 10^{7} M\rm M_{\odot}$, the constant-density models overpredict this fraction, since the stellar density tends to decrease with decreasing black hole mass. For black holes with mass larger than $M_{b}$, the constant-density models underpredict the fraction of mass from stellar accretion.
The predicted mass evolution in Figures 3 and 4 illustrates that conditions for black hole growth vary widely among galaxies. Of the 30 sources we consider here, 20% (6) are predicted to have at least doubled the black hole mass in the past gigayear, while in 8 galaxies, stellar accretion from binary break-ups accounts for less than 10% of the black hole mass.
The evolution calculations highlight two galaxies that exhibit exceptional SMBH growth rates over gigayear timescales: NGC 4342 and NGC 4026. The central SMBH in NGC 4342, a lenticular (S0) galaxy, in our calculations has formally acquired over 99% of its mass from the binary-capture mechanism over the last gigayear, and is predicted to grow by another order of magnitude over the next 5 Gyr. However, our predictions likely overestimate the potential of the stellar accretion growth channel. First, the measured velocity dispersion is significantly higher than the 100 km/s value in our model, which would mean we overestimate the binary break-up rate because our gravitational focusing factor is too high. Cretton & van den Bosch (1999) report $v_{\rm disp}\approx 400$ km/s from the Hubble Space Telescope. Second, the binary fraction in this early-type galaxy, presumably composed of lower-mass, older stars, is likely lower than our fiducial value of 0.5. Adopting conservative values of $v_{\rm disp}=400$ km/s and $f_{\rm bin}=0.1$ (see Basri & Reiners, 2006, for example), our model still predicts that over 20% of the SMBH mass in NCG 4342 came from this growth channel in the last gigayear, and more than 95% over the past 5 Gyr.
Similarly, the SMBH in NGC 4026, also an early-type S0 galaxy, evolves from an initial mass of $1.82\times 10^{8}$ $\rm M_{\odot}$ to over $10^{9}\,M_{\odot}$ within a couple of gigayears in our model. The fraction of mass from binary capture is formally over 95% (Fig. 4). Like the case of NGC 4342, this galaxy has a measured velocity dispersion that is greater than our fiducial value. Adopting a velocity disperson of 250 km/s (Gültekin et al., 2009), and the lower value of $f_{\rm bin}=0.1$, we find that binary captures have contributed to about 20% of the present-day SMBH mass in the past gigayear, and over 80% of its mass in the past 5 Gyr.
We conclude that for the population of galaxies considered here, the binary-capture channel can play a significant, and possibly even dominant role in black hole growth.
4 Closer to home: the case of LMC*
Han et al. (2025) recently proposed the existence of a $6\times 10^{5}$ $\rm M_{\odot}$ black hole within the Large Magellanic Cloud (LMC). Their inference is based on the locations and velocities of the observed HVSs from the HVS Survey (Brown et al., 2012). This analysis incorporated Gaia Data Release 3 astrometry (Gaia Collaboration et al., 2023) to reveal that half of the HVS Survey sources originated in the LMC. With forward modeling using the ejection speed parameterization from Bromley et al. (2006) and an HVS production rate of $k=2$ Myr-1, Han et al. (2025) reproduced the observed phase-space distribution of these HVSs, including the velocity boost from LMC motion in the Galactic frame and the observed clustering on the sky within the context of the survey selection. The success of this modeling requires the presence of LMC.
The results in Han et al. (2025) have implications for the growth of the dwarf galaxy’s massive black hole. The observed cluster of nine main-sequence, B-type HVSs linked to an LMC origin, found in the constellation Leo, turn out to be the tip of an iceberg representing binary star encounters with LMC. First, these stars are at the edge of the HVS Survey window, which do not include a large swath of sky around the LMC where the majority of HVSs are expected to reside. From the reported rate of HVS production required to match observations, these nine stars are associated with $\sim$800 other ejected stars with similar masses that lie beyond the surveyed sky. Second, B-stars are comparatively rare, and are the tip of another iceberg from a stellar population perspective. With a Salpeter initial mass function (IMF), stars with mass $m\geq 2$ $\rm M_{\odot}$ — the range considered by Han et al. (2025) — represent less than 2% of all stars with masses greater than 0.1 $\rm M_{\odot}$. The nine HVSs from LMC thus suggest a binary disruption rate greater than $10^{-4}$ yr-1. With an average mass around $0.3$ $\rm M_{\odot}$ captured onto tight orbit around LMC per disruption, the SMBH can acquire a significant fraction of its mass over a few gigayears.
There are significant uncertainties in this growth rate. A major source is the initial mass function, since the HVS Survey only sampled the most massive stars. A top-heavy IMF would reduce the rate we infer here — the observed stars then are the tip of a less significant iceberg. Other sources of uncertainty include the details that affect the Hills mechanism and its outcome, including the distribution of binary orbits, the periastron distribution (empty or full loss cone Lightman & Shapiro, 1977; Stone & Metzger, 2016), and the mass distribution around the SMBH that impacts the speed (and interpretation) of the HVS candidates (e.g., Bromley et al., 2006; Sari et al., 2010; Rossi et al., 2014; Han et al., 2025).
A check on our overall approach (§2.2) is a comparison between our predicted binary break-up rate and the Han et al. (2025) LMC HVS study. With a velocity dispersion of 35 km/s, characteristic of the central region of the LMC (Vasiliev, 2018), and (in absence of a compelling value from the literature) a central density of $\rho_{0}\approx 750$ $\rm M_{\odot}$/pc3 from Equation (4), we find $k=2.0\times 10^{-4}$, similar to the estimate based on the HVS number count reported in Han et al. (2025). Our model predicts that about 10% of LMC’s mass came from stellar accretion over the last gigayear, and 30% in the past 5 Gyr. Although binary accretion on its own cannot account for the emergence of a $10^{6}$ /msun/ black hole, the mechanism can help build a SMBH from a smaller seed object. A recent detailed study of black hole growth in globular clusters came to a similar conclusion (González Prieto et al., 2025).
5 Conclusion
This analysis explores SMBH growth from accretion of stars captured through binary break-up. We presented a simple theoretical framework to assess the effectiveness of this growth channel (§2). We then applied it to 30 galaxies hosting SMBHs with well-constrained masses, predicting both past and future growth by this mechanism (§3). In most cases, this growth channel can account for at least 10% of the black hole mass acquired over the past Gigayear. In 20% of the galaxies, 50% or more of their mass was obtained in this way.
Our model, an extension of one proposed by Bromley et al. (2012), has its limitations. These include uncertainties in the demographics and intrinsic properties of binary stars in the central regions of galaxies, including whether the orbital distribution of binaries is consistent with a full loss cone (Lightman & Shapiro, 1977), as we have assumed here. A comparison with data from HVSs that originate from interaction with the LMC’s black hole (§4) is a promising indicator that some of these uncertainties may be reduced in the future. Broader comparisons with tidal disruption events and their rate of occurrence (e.g., Stone & Metzger, 2016) may also inform our model.
The scenarios described here, with steady-state rates of binary star-SMBH encounters, may be extended to include binary captures during merger events of dwarf galaxies and even major mergers between comparably sized galaxies. Our preliminary explorations indicate that such stochastic events might supply a wealth of new stellar binaries into the central regions of galaxies. This possibility may further enhance the potential of the Hills mechanism not only as a signpost of supermassive black holes, but as a way to grow them.
Acknowledgments
We thank C. Hannah for generously providing the SMBH data that formed the foundation of this analysis, as well as guidance from A. Seth. AJ is grateful for support from the University of Utah’s Undergraduate Research Opportunity Program.